# AT 2025abao: the fourth luminous red nova in M 31

### v2

@@ 4 Spectral evolution
On 2025 November 3 (phase d), we obtained a high-resolution () Echelle spectrum with the Copernico 1.82-m telescope, in which we can identify specific spectral features. The main spectral features identified in this spectrum are highlighted in Fig. 12. H and H have a narrow P Cygni absorption profile on top of a much broader component, with a velocity of 40 km s-1 (Figure 12, right panel). Along with the H lines, we clearly identify prominent Na i D doublet features, all the components well discerned. However, while the absorption lines of the Na i D doublet at redshift zero are unresolved and hence can be attributed to material along the line of sight within the Milky Way, the Na i D doublet lines at the redshift of M 31 are resolved, and have a FWHM of around 20-25 km s-1 (see Fig. 18 in Appendix A, top panel). This FWHM velocity is similar to those exhibited by other metal lines visible already in the d spectrum (see below), hence we consider these Na i D features as intrinsic to the transient. A prominent absorption feature is observed at 5890.9 Å, which is unlikely to be an additional Na I component, and whose nature will be discussed in Appendix A. Lines of the Mg I triplet at 5160-5180 Å are also detected, with a FWHM velocity similar to that of the Na i D lines. More metal lines are detected bluewards of 5200 Å, including from Fe ii, which also have a FWHM velocity of 25 km s-1. Finally, broad Ca ii H&K absorption lines are visible in the UV region, with a much higher FWHM velocity of km s-1 (Figure 12, left panel), due to its opaqueness.
@@
In the spectra taken at about the -band maximum light, together with a double-horned H line, a faint He i 6678 is the only other transition seen in emission, on top of a continuum peaking at 4800-4900 Å (indicating 6000 K, consistent with the SED inferred in Sect. 3.6). Most notable features are the deep and strong absorptions of the Ca ii H&K UV doublet, Na i D , and a forest of metal lines. In particular, the Fe ii multiplet 42 ( 4924, 5018, 5169) transitions are clearly detectable.
@@
Already by the d spectrum, the continuum temperature has declined to 5200 K (with the spectral continuum emission peaking at around 5600 Å). The gas cooling is confirmed by the disappearance of the He I 6678 line. Otherwise, the spectrum did not change significantly, with the most prominent metal transitions still clearly present. In the d spectrum, the continuum peaks at around 5900 Å, indicating a temperature of only 4900 K. Emission lines, including H, have completely disappeared, and visible features are only in absorption. Multiple metal lines are identified in the bluer region, together with the usual Ca ii (both H&K and IR triplet) and Na i D . Little has changed in the spectrum at d. A broad basin is visible between 4800 and 5300 Å, a feature commonly seen in the spectra of Sun-like stars, and attributed to the absorption due to the CH radical molecule.
@@
In the d spectrum, as already in the d spectrum, the continuum peak has shifted to around 6800 Å (4250 K), maintaining a steep decline towards the UV likely due to line blanketing, while being flat at redder wavelengths. The strongest features are now the absorptions from the Ca ii NIR triplet, though the Ca ii H&K absorption lines are also recognizable, together with the depression due to CH, and even more numerous metal lines. Balmer lines are now weak or absent.
@@
On December 10 (phase d), we took another Echelle spectrum, which reveals an astonishing peculiarity: the H line presents a counter P Cygni profile, i.e. a narrow absorption red-wards of the centre of the emission line. The velocity of the minimum of this feature indicates an infalling velocity of km s-1. The component is confirmed as it is detected also in the H line (see Fig. 13, where its profile is overlapped with that of H). This second Echelle spectrum also shows a crowded line population in the 5850-5900 Å region: a blueshifted He i 5876 emission line, two Na i D doublets (one due to line-of-sight material in the MW, and the other intrinsic to AT 2025abao), and two more lines at 5890.9 and 5896.9, also discussed in Appendix A.
@@
In the d spectrum of AT 2025abao, the first signs of molecular absorption bands: TiO features start to appear between 6000 and 8000 Å, in the form of two sawtooth drops in the continuum, sharp on the blue edge and then gradually rising towards the red edge. This trend of molecules formation is confirmed in the d spectrum. Molecular bands from VO and TiO are clearly present in the regions 6100-6700 Å and 7100-7800 Å, together with a rise at the start of the NIR, in the range 8400-8800 Å, also compatible with a molecular band. The continuum peaks at around 7800 Å (3700 K), with a temperature trend consistent with that expected for a moderately faint LRN. Balmer lines appear to be strong again in the and d spectra, while the continuum is now red and its temperature has decreased to 3600 and 3500 K, respectively.
@@ 4.1 Spectral line identification
On November 25 ( d), four medium-resolution () spectra were obtained with 10.4m GTC+OSIRIS and the R2500-class grisms, covering wavelengths from the to the bands. In these spectra, we identified the most prominent metal lines, highlighting that these features are real and not noise patterns, since for each grism we took four exposures in sequence and checked whether a line was present in all individual spectra.
In the AT 2025abao spectrum, H shows a narrow P Cygni profile with a dominant absorption with a minimum blue-shifted by km s-1, and a weaker and narrow emission component. However, on the red side, a sharper counter P Cygni is also present. This feature is common in LRN spectra (see AT 2011kp; Pastorello et al. 2019), and can be interpreted as outflowing material from the central system through the second Lagrangian point (Pejcha et al., 2016a) in a direction opposite to the observer. Other Balmer lines are detected (H, H, and H, and possibly also H), dominated by the emission component. Ca ii H&K are relatively broad (FWHM km s-1) and in pure absorption. Other notable absorption features are the Ca ii NIR triplet, and the Na i D doublet. All other features, while present and identified, are less strong.
In Fig. 14, we compare the full GTC+OSIRIS mid-resolution spectrum of AT 2025abao with that of AT 2011kp during late phases ( d) from Pastorello et al. (2019), who also performed a similar line identification in a mid-resolution spectrum. The spectrum of AT 2025abao was taken 20 days after maximum, while that of AT 2011kp was obtained after more than three months. While the spectral continuum shows some differences, the most important metal lines are identified in both spectra. This is an indication of a much faster evolution of AT 2025abao. Thus, a similar set of lines is identified and visualized in both spectra: multiplets from Fe ii, Fe i, Mg i, Sc ii, Ba ii (the strongest line being 6497) and Ti ii are retrieved. Single lines from H, O i, Ca ii, and Na I are also found. In the red extreme, in a region not strongly contaminated by telluric absorption bands, we also identify the Na I 8183,8195 doublet, O i 7774 and 8446. Many lines, not marked in Fig. 14, can be attributed to a multitude of transitions from Ti ii, as suggested also by Pastorello et al. (2019).
@@ 5 Discussion and conclusion
AT 2025abao is the most recent example of a LRN in M 31, and the fourth in the last 40 years in this galaxy, with an observed LRN events rate of . All of them have peak absolute magnitudes in the range to mag. Brighter objects would have been easily detected, while fainter ones could be lurking among the Novae population. For comparison, only one object (V838 Mon) was discovered with this luminosity range in a similar lapse of time in the MW. Based on the estimations of Kochanek et al. (2014), we calculate an event rate of just 0.04 yr-1 within the same magnitude range (0.07 yr-1 of events brighter than mag minus 0.03 yr-1 brighter than mag). Therefore, it seems that this kind of transients is more than twice as common in the Andromeda Galaxy than in the MW. The stellar mass of M 31 has been estimated to be (Barmby et al., 2006; Tamm et al., 2012; Kafle et al., 2018), while the equivalent for the Milky Way is (Licquia and Newman, 2015; Bland-Hawthorn and Gerhard, 2016). Hence, the Andromeda Galaxy is times more massive in stellar content and likely in number of stars. If we scale up the Galactic LRN rate in the to mag range from Kochanek et al. (2014) by this ratio, we find a rate that is very similar to the observed one. The LRN population is dominated by faint events, with 86 % of them having between and mag, while those in the to mag range make only 8%, and the more luminous than mag just 6% (however, Karambelkar et al. 2023 predict an even lower rate for these brighter events). Therefore, while we expect more than one LRN per year in M 31 ( for mag according to Kochanek et al. 2014, times equals ), the rate of events should only be (or less according to Karambelkar et al. 2023), making them particularly rare.
@@
Usually, the progenitors of LRNe are found to be yellow stars, often supergiants in the cases of the most luminous events (Mauerhan et al., 2015; Blagorodnova et al., 2017; Pastorello et al., 2021b; Blagorodnova et al., 2021; Cai et al., 2022a). However, the association of AT 2025abao with WNTR23bzdiq, a slowly variable AGB star (K25), makes it an exception. K25 interpreted this object as a common-envelope event and predicted a merging event as the outcome. AGB stars are so inflated that the outer layers of the atmosphere can be considered like a hot vacuum. Therefore, the companion star - especially if the mass ratio is particularly large - could pass through it, remaining engulfed within the envelope and without significantly changing the global colour of the system, which would remain red, as observed by K25.
@@
We constructed the pre-outburst SED of AT 2025abao at the time of the last local peak in the light curve (at 333 d), before the optically thick phase. We used only the -bands data from ZTF, and since there was no contemporary -band imaging, we assumed a constant colour from the only pre-outburst -band public observation by ZTF (at 411 d, mag). Subsequently, we corrected the SED for the same Galactic extinction adopted by K25 and accounting for their dust parameters obtained with the radiative transfer code DUSTY (Ivezic and Elitzur, 1997) for the quiescent progenitor (i.e. , K, K, , , silicate composition). These parameters, according to our DUSTY simulation, result in an additional circumstellar extinction of mag. We compared this corrected SED with the ATLAS9 stellar atmospheric models by Castelli and Kurucz (2003)151515retrieved here: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/grid/ck04models/. We adopted the models with and log. We find a good match with a model with a photospheric temperature of 4000 K, and a stellar radius of about 540 . The match of the ATLAS9 models to the precursor SED is shown in Fig. 15. These parameters are consistent with those of an orange supergiant, but also with an expanded AGB star, though hotter and more luminous (log vs. 4.20) than the progenitor parameters obtained by K25. If we instead adopt our distance and reddening estimates (Sect. 2), the corrected SED is fitted with a model having K, a radius of , and a corresponding log, still hotter and more luminous than the quiescent progenitor.
@@
We also used DUSTY to model the SED of the precursor of AT 2025abao observed by SPHEREx a few months before the outburst, at a phase close to the optical minimum. Since the broad molecular absorptions at 1.4-1.5 and 1.8-2.2 m are large deviations from a BB, we opted for an atmospheric model of a cool star as the radiation source. We adopted the PHOENIX models161616retrieved here: https://pollux.oreme.org/explore/BT-Dusty/ (Husser et al., 2013) with effective temperatures between 2500 and 3500 K, , and log. Good matches are found for the model with K and , which is higher than that found for the quiescent progenitor. The matches of the PHOENIX+DUSTY models to the SED from SPHEREx are shown in Fig. 16. The inner radius of the shell of dust is ( AU), while the radius of the source is 350 , the same as the quiescent progenitor. However, the dust mass we obtain, following the procedures of Reguitti et al. (2026), is only , one order of magnitude lower than K25. This is due to the inner radius of the dust shell being much smaller.
@@
K25 estimated a mass of the progenitor of based on evolutionary paths from MESA isochrones and stellar tracks. Using the empirical relation between absolute magnitude and the mass of the progenitor by Cai et al. (2022a) for an object with mag during the plateau phase171717Obviously, AT 2025abao does not exhibit a genuine secondary red peak, but instead a long plateau. However, the relation of Cai et al. (2022a) can still be used, considering the -band absolute magnitude at the end of the plateau. Consequently, in Appendix B, the Cai et al. (2022a) relation has been updated using also the AT 2025abao data., we obtain , in excellent agreement with their estimation. Based on the peak luminosity and similarities in the spectral evolution, we can also state that the progenitor mass of AT 2025abao must be similar to that of V838 Mon, which was (Pastorello et al., 2023; Tylenda, 2005) and slightly more massive than that of AT 2019zhd (3.4 for Pastorello et al. 2021a; 6 for Chen and Ivanova 2024).
To provide more physical parameters, we considered the analytical expressions derived by Matsumoto and Metzger (2022). In particular, their Equation 16 allows to estimate the mass of the ejected material by measuring the duration of the plateau and the bulk velocity of the ejecta . For the plateau length, we assume days, as in the -band it started one week before maximum light, and ended at d. For the ejecta velocity, we adopted the FWHM of the Hlines during the plateau (as in Cai et al. 2022a), precisely in the and d spectra, which give km s-1. Under these assumptions, we calculate and ejecta mass of 0.63 .
@@
AT 2025abao is a LRN exhibiting a plateau instead of a double-peaked light curve. Its progenitor has been established to be an early AGB (K25), hence an expanded star with an H-rich envelope. The photometric behaviour of AT 2025abao, similar to that of a Type IIP SN, can be explained with the recombination of a large quantity of H, previously ionized by the shocks following the merger of the two stars (MacLeod et al., 2017). Therefore, we propose that the observed dichotomy in LRNe – with some showing two clear light curve peaks (the first blue and the second red) and others with a plateau instead – is linked to the extent of the common envelope and the amount of H. Hence, systems having compact progenitors or with low H content would be responsible for double-peaked LRNe, whilst events with extended progenitors and massive outer H layers would produce LRNe with flat light curves (see also Howitt et al. 2020; Twum et al. 2026). The total mass of the system instead influences the luminosity of the event (as established by Kochanek et al. 2014; Blagorodnova et al. 2021; Cai et al. 2022a). Another possibility is that the plateau luminosity is maintained long by shocks produced by the collision of shells, together with the H-recombination (Matsumoto and Metzger, 2022; Kirilov et al., 2025).
@@
The narrow absorption visible on top of the H profile in the spectra at maximum light (Fig. 11, right panel), with the red ‘horn’ fainter than the blue one, can be explained with a fast outflow far from being spherically symmetric (though it is still possible it could be axisymmetrical). We propose a scenario in which an excretion circumbinary disk (which may form if not all of the envelope is ejected; Ivanova et al., 2013b) or a torus-shaped structure surrounds the system. It would be observed at a nearly edge-on viewing angle, and would obscure the far receding side of the CSM, making the red-shifted portion of it more attenuated and thus fainter (see also Kirilov et al., 2025). We present a sketch of this scenario in Fig. 17.
@@
In conclusion, it will be crucial to observe AT 2025abao years after the event with space telescopes, especially in the infrared domain (e.g. with JWST, as done by Karambelkar et al. 2025a, or with the forthcoming Roman, Akeson et al., 2019), to see what kind of outcome remains, and if the system returns to a state similar to the pre-LRN one, with an inflated red star. We note that a red (super)giant survivor was observed after the coalescence of a number of LRNe (Steinmetz et al., 2025; Reguitti et al., 2026). Future observations will also constrain the variability of the post-merger remnant, which may confirm whether the stellar components of the system had completely coalesced. In fact, very late-time observations of V838 Mon by Goranskij et al. (2020) revealed a slow brightening in the optical, which raises doubts if a complete merger really occurred in that system.
@@
The observed calibrated magnitudes are tabulated in the photometry tables, which are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. All the spectra are released on the WISeREP interface (https://www.wiserep.org/).

### v3

@@ 4.1 Main spectral evolution
On 2025 November 3 (phase d), we obtained a high-resolution () Echelle spectrum with the Copernico 1.82-m telescope, in which we can identify specific spectral features. The main spectral features identified in this spectrum are highlighted in Fig. 12. H and H have a narrow absorption profile with a width of 45 km s-1, on top of a much broader emission component (450 km s-1, Fig. 12, right panel). Along with the H lines, we clearly identify prominent Na i D doublet features, and all the components well discerned. However, while the absorption lines of the Na i D doublet at redshift zero are unresolved and hence can be attributed to material along the line of sight within the MW, the Na i D doublet lines at the redshift of M 31 are resolved and have a FWHM of around 20-25 km s-1 (see Fig. 18 in Appendix A, top panel). This FWHM is similar to those exhibited by other metal lines visible in the d spectrum (see below); hence, we consider these Na i D features as intrinsic to the transient. A prominent absorption feature is observed at 5890.9 Å, which is unlikely to be an additional Na I component, and whose nature is discussed in Appendix A. Lines of the Mg I triplet at 5160–5180 Å are also detected, with a FWHM similar to that of the Na i D lines. More metal lines are detected bluewards of 5200 Å, including from Fe ii, which also have a FWHM of 25 km s-1. Finally, broad Ca ii H&K absorption lines are visible in the UV region, with a much higher FWHM of km s-1 (Fig. 12, left panel). The broadness of the Ca ii H&K lines is due to the elevated opacity of those transition lines, as effects of saturation are noticeable and possibly also blending with other metal lines. However, it is still possible to have a fast-moving outflow, produced after the merger of the stars’ cores, perpendicular to the orbital plane (see Metzger and Pejcha 2017; Gagnier and Pejcha 2025) and with a higher velocity than the material along the equatorial disc.
@@
In the spectra taken at about the -band maximum light, together with a double-horned H line, a faint He i 5876 is the only other transition seen in emission, on top of a continuum peaking at 4800–4900 Å (indicating 6000 K, which is consistent with the SED inferred in Sect. 3.6). The most notable features are the deep and strong absorptions of the Ca ii H&K UV doublet, Na i D, and a forest of metal lines. In particular, the Fe ii multiplet 42 ( 4924, 5018, 5169) transitions are clearly detectable.
@@
By the d spectrum, the continuum temperature has already declined to 5200 K (with the spectral continuum emission peaking at around 5600 Å). The gas cooling is confirmed by the disappearance of the He I 6678 line. Otherwise, the spectrum did not change significantly, with the most prominent metal transitions still clearly present. In the d spectrum, the continuum peaks at around 5900 Å, indicating a temperature of only 4900 K. Emission lines, including H, have completely disappeared, and visible features are only in absorption. Multiple metal lines are identified in the bluer region, together with the usual Ca ii (both H&K and IR triplet) and Na i D. Little has changed in the spectrum at d. A broad basin is visible between 4800 and 5300 Å, a feature commonly seen in the spectra of Sun-like stars, and attributed to the absorption due to the CH radical molecule.
@@
In the d spectrum, as already in the d spectrum, the continuum peak has shifted to around 6800 Å (4250 K), maintaining a steep decline towards the UV likely due to line blanketing, while being flat at redder wavelengths. The strongest features are now the absorptions from the Ca ii NIR triplet, though the Ca ii H&K absorption lines are also recognisable, together with the depression due to CH, and even more numerous metal lines. Balmer lines are now weak or absent.
@@
On December 10 (phase d), we took another Echelle spectrum, which reveals a peculiarity: the H line presents a narrow absorption redwards of the centre of the emission. The velocity of the minimum of this feature relative to the emission peak is km s-1. The same structure is also detected in the H line (see Fig. 13, where its profile is overlapped with that of H). The feature looks similar to a counter P Cygni profile. An inverse P Cygni profile has been observed before in LRN spectra (see AT 2011kp; Pastorello et al. 2019) and in another IGT, AT 2008jd (Berger et al., 2009). In the latter case, it was interpreted as infalling material from a disc on the orbital plane towards the central system through the second Lagrangian point (see also Pejcha et al., 2016a). However, in the case of AT 2025abao the emission and absorption components might not be correlated and be dynamically distinct (Mason and Shore, 2022). This second Echelle spectrum also shows a crowded line population in the 5850-5900 Å region: a blueshifted He i 5876 emission line, two Na i D doublets (one due to line-of-sight material in the MW, and the other intrinsic to AT 2025abao), and two more lines at 5890.9 and 5896.9, also discussed in Appendix A.
@@
In the d spectrum of AT 2025abao, the first signs of molecular absorption bands are seen: TiO features start to appear between 6000 and 8000 Å, in the form of two sawtooth drops in the continuum, which are sharp on the blue edge and then gradually rise towards the red edge. This trend of molecule formation is confirmed in the d spectrum. Molecular bands from VO and TiO are clearly present in the regions 6100–6700 Å and 7100–7800 Å, together with a rise at the start of the NIR, in the 8400–8800 Å range, also compatible with a molecular band. The continuum peaks at around 7800 Å (3700 K), with a temperature trend consistent with that expected for a moderately faint LRN. Balmer lines appear to be strong again in the and d spectra, while the continuum is now red and its temperature has decreased to 3600 and 3500 K, respectively.
@@ 4.2 Spectral line identification
On November 25 ( d), four medium-resolution () spectra of AT 2025abao were obtained with 10.4m GTC+OSIRIS and the R2500-class grisms, covering wavelengths from the to the bands. In these spectra, we identified the most prominent metal lines, highlighting that these features are real and not noise patterns, since for each grism we took four exposures in sequence and checked whether a line was present in all individual spectra.
In the GTC spectrum, H shows a narrow P-Cygni profile with a dominant absorption with a minimum blueshifted by km s-1 (although the blue wing extends down to km s-1), and a weaker and narrow emission component. However, on the red side, a sharper counter P Cygni is also present. Other Balmer lines are detected (H, H, H, and possibly H), dominated by the emission component. Ca ii H&K are relatively broad (FWHM km s-1) and in pure absorption. Other notable absorption features are the Ca ii NIR triplet, and the Na i D doublet. All other features, while present and identified, are less strong.
In Fig. 14, we compare the full GTC+OSIRIS mid-resolution spectrum of AT 2025abao with that of AT 2011kp during late phases ( d) from Pastorello et al. (2019), the authors of which also performed a similar line identification in a mid-resolution spectrum. The spectrum of AT 2025abao was taken 20 days after maximum, while that of AT 2011kp was obtained after more than three months. While the spectral continuum shows some differences, the most important metal lines are identified in both spectra. This is an indication of a much faster evolution of AT 2025abao. Thus, a similar set of lines is identified and seen in both spectra: multiplets from Fe ii, Fe i, Mg i, Sc ii, Ba ii (the strongest line being 6497), and Ti ii are retrieved. Single lines from H, O i, Ca ii, and Na I are also found. In the red extreme, in a region not strongly contaminated by telluric absorption bands, we also identify the Na I 8183,8195 doublet, O i 7774 and 8446. Many lines, not marked in Fig. 14, can be attributed to a multitude of transitions from Ti ii, as also suggested by Pastorello et al. (2019).
@@ 5 Discussion and conclusion
AT 2025abao is the most recent example of an LRN in M 31, and the fourth in the last 40 years in this galaxy, with an observed LRN-event rate of . All of them have peak absolute magnitudes in the to mag range. Brighter objects would have been easily detected, while fainter ones could be lurking among the Novae population. For comparison, only one object (V838 Mon) was discovered with this luminosity range in a similar lapse of time in the MW. Based on the estimations of Kochanek et al. (2014), we calculated an event rate of just 0.04 yr-1 within the same magnitude range (0.07 yr-1 of events brighter than mag minus 0.03 yr-1 brighter than mag). Therefore, it seems that this kind of transient is more than twice as common in the Andromeda galaxy than in the MW. The stellar mass of M 31 has been estimated to be (Barmby et al., 2006; Tamm et al., 2012; Kafle et al., 2018), while the equivalent for the MW is (Licquia and Newman, 2015; Bland-Hawthorn and Gerhard, 2016). Hence, the Andromeda galaxy is times more massive in terms of stellar content and, likely, in its number of stars. If we scale up the Galactic LRN rate in the to mag range from Kochanek et al. (2014) by this ratio, we find a rate that is very similar to the observed one. The LRN population is dominated by faint events, with 86 % of them having between and mag, while those in the to mag range make up only 8%, and those more luminous than mag just 6% (however, Karambelkar et al. 2023 predicted an even lower rate for these brighter events). Therefore, while we expect more than one LRN per year in M 31 ( for mag according to Kochanek et al. 2014, times equals ), the rate of events should only be (or less according to Karambelkar et al. 2023), making them particularly rare.
@@
Usually, the progenitors of LRNe are found to be yellow stars, often supergiants in the cases of the most luminous events (Mauerhan et al., 2015; Blagorodnova et al., 2017; Pastorello et al., 2021b; Blagorodnova et al., 2021; Cai et al., 2022a). However, the association of AT 2025abao with WNTR23bzdiq, a slowly variable AGB star (K25), makes it an exception. K25 interpreted this object as a common-envelope event and predicted a merging event as the outcome. AGB stars are so inflated that the outer layers of the atmosphere can be considered a hot vacuum. Therefore, the companion star –especially if the mass ratio is particularly large– could pass through it, remaining engulfed within the envelope without significantly changing the global colour of the system, which would remain red, as observed by K25.
@@
We constructed the pre-outburst SED of AT 2025abao at the time of the last local peak in the light curve (at 333 d), before the optically thick phase. We only used the -band data from ZTF, and since there was no contemporary -band imaging, we assumed a constant colour from the only pre-outburst -band public observation by ZTF (at 411 d, mag). Subsequently, we corrected the SED for the same Galactic extinction adopted by K25 and accounting for their dust parameters obtained with the radiative transfer code DUSTY (Ivezic and Elitzur, 1997) for the quiescent progenitor (i.e. , K, K, , , silicate composition). These parameters, according to our DUSTY simulation, result in an additional circumstellar extinction of mag. We compared this corrected SED with the ATLAS9 stellar atmospheric models by Castelli and Kurucz (2003)151515Retrieved here: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/grid/ck04models/. We adopted the models with and log. We find a good match with a model with a photospheric temperature of 4000 K, and a stellar radius of about 540 . The match of the ATLAS9 models to the precursor SED is shown in Fig. 15. These parameters are consistent with those of an orange supergiant, but also with an expanded AGB star, though hotter and more luminous (log vs. 4.20) than the progenitor parameters obtained by K25. If we instead adopt our distance and reddening estimates (Sect. 2), the corrected SED is fitted with a model that has K, a radius of , and a corresponding log, which is still hotter and more luminous than the quiescent progenitor.
@@
We also used DUSTY to model the SED of the precursor of AT 2025abao observed by SPHEREx a few months before the outburst, at a phase close to the optical minimum. Since the broad molecular absorptions at 1.4-1.5 and 1.8-2.2 m are large deviations from a BB, we opted for an atmospheric model of a cool star as the radiation source. We adopted the PHOENIX models161616Retrieved here: https://pollux.oreme.org/explore/BT-Dusty/ (Husser et al., 2013) with effective temperatures, , between 2500 and 3500 K, , and log. Good matches are found for the model with K and , which is higher than that found for the quiescent progenitor. The matches of the PHOENIX+DUSTY models to the SED from SPHEREx are shown in Fig. 16. The inner radius of the shell of dust is ( AU), while the radius of the source is 350 , the same as the quiescent progenitor. However, the dust mass we obtain, following the procedures of Reguitti et al. (2026), is only ; that is, one order of magnitude lower than K25. This is due to the inner radius of the dust shell being much smaller.
@@
K25 estimated a mass of the progenitor of based on evolutionary paths from MESA isochrones and stellar tracks. Using the empirical relation between absolute magnitude and the mass of the progenitor by Cai et al. (2022a) for an object with mag during the plateau phase171717Obviously, AT 2025abao does not exhibit a genuine secondary red peak, but instead a long plateau. However, the relation of Cai et al. (2022a) can still be used, considering the -band absolute magnitude at the end of the plateau. Consequently, in Appendix B, the Cai et al. (2022a) relation has been updated using also the AT 2025abao data., we obtained , which is in excellent agreement with their estimation. Based on the peak luminosity and similarities in the spectral evolution, we can also state that the progenitor mass of AT 2025abao must be similar to that of V838 Mon, which was (Pastorello et al., 2023; Tylenda, 2005) and slightly more massive than that of AT 2019zhd (3.4 for Pastorello et al. 2021a; 6 for Chen and Ivanova 2024).
To infer additional physical parameters, we considered the analytical expressions derived by Matsumoto and Metzger (2022). In particular, their Equation 16 allowed us to estimate the mass of the ejected material, , by measuring the duration of the plateau, , and the bulk velocity of the ejecta, . For the plateau length, we adopted a value of days, as it started in the band one week before maximum light and ended at d. For the ejecta velocity, we adopted the FWHM of the H line estimated during the plateau (as in Cai et al. 2022a), precisely in the and d spectra; hence, km s-1. The same velocity was obtained by measuring the FWHM of the broad component in the H emission line in the first Echelle spectrum, and from the blue wing of the P Cygni profile of the same line (and H) in the GTC spectrum. Under these assumptions, we calculated an ejecta mass of 0.63 . This value roughly coincides with the one found by Wu et al. (2026), who modelled our bolometric light curve. The associated kinetic energy is erg; as a comparison, the total kinetic+radiated energy emitted by V838 Mon was erg (Tylenda and Soker, 2006). The collision of the two stellar cores deposits sufficient energy (see also Bally and Zinnecker, 2005) in the outer layers to lift the ejecta at that encountered high speed.
@@
AT 2025abao is an LRN exhibiting a plateau instead of a double-peaked light curve. Its progenitor has been established as an early AGB (K25), that is, an expanded star with an H-rich envelope. The photometric behaviour of AT 2025abao, similar to that of a Type IIP SN, can be explained with the recombination of a large quantity of H, previously ionised by the shocks following the merger of the two stars (MacLeod et al., 2017). Therefore, we propose that the observed dichotomy in LRNe –with some showing two clear light-curve peaks (the first blue and the second red) and others with a plateau instead– is linked to the extent of the common envelope and the amount of H. Hence, systems with compact progenitors or a low H content would be responsible for double-peaked LRNe, whilst events with extended progenitors and massive outer H layers would produce LRNe with flat light curves (see also Howitt et al. 2020; Twum et al. 2026). The total mass of the system instead influences the luminosity of the event (as established by Kochanek et al. 2014; Blagorodnova et al. 2021; Cai et al. 2022a). Another possibility is that the plateau luminosity is maintained long by shocks produced by the collision of shells, together with the H recombination (Matsumoto and Metzger, 2022; Kirilov et al., 2025).
@@
The double-peaked H profile in the spectra at maximum light (Fig. 11, right panel), with the red ‘horn’ fainter than the blue one, can be explained by an excretion circumbinary disc (which may form if not all of the envelope is ejected; Ivanova et al., 2013b) or a torus-shaped structure surrounds the system. We propose a scenario in which this disc or torus would be observed at a nearly edge-on viewing angle and would obscure the receding side of the CSM, making the redshifted portion of it more attenuated and thus fainter (see also Kirilov et al., 2025). A fast outflow (450 km s-1), far from being spherically symmetric (though it is still possible it could be axisymmetrical), is instead responsible for the broad emission component underneath the double-peaked profile. This model is similar to that presented by Kashi and Soker (2017). We present a sketch of this scenario in Fig. 17.
| Mass of the primary star () | |
| Radius of the primary star () | |
| Temperature of the primary star () | 3100 K |
| Viewing angle | |
| Ejecta mass () | |
| Fast outflow velocity () | km s-1 |
| Kinetic energy () | erg |
| Velocity of the material on the orbital plane | km s-1 |
| Total radiated energy | erg |
@@
In conclusion, it will be crucial to observe AT 2025abao years after the event with space telescopes, especially in the infrared domain (as done by Karambelkar et al. 2026, or with the forthcoming Roman, Akeson et al., 2019), to see what kind of outcome remains and if the system returns to a state similar to the pre-LRN one, with an inflated red star. We note that a red (super)giant survivor was observed after the coalescence of a number of LRNe (Steinmetz et al., 2025; Reguitti et al., 2026). Future observations will also constrain the variability of the post-merger remnant, which may confirm whether the stellar components of the system had completely coalesced. In fact, very-late-time observations of V838 Mon by Goranskij et al. (2020) revealed a slow brightening in the optical, which raises doubts as to whether a complete merger really occurred in that system.
@@
The observed calibrated magnitudes are tabulated in the photometry table, which is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. All the spectra are released on the WISeREP interface (https://www.wiserep.org/).